OpenStudy (christinestutes):

Graph the equation 2x-4 =0

OpenStudy (christinestutes):

Do I find x and y intercepts?

jimthompson5910 (jim_thompson5910):

y is not in this equation at all, right?

OpenStudy (christinestutes):

or y=mx+b 0= 2x-4 4=2x x=2 (0,2) y intercept is 2?

OpenStudy (christinestutes):

right Jim

OpenStudy (christinestutes):

Y is not in equation which is what is confusing me

jimthompson5910 (jim_thompson5910):

ok then just solve for x. What do you get?

jimthompson5910 (jim_thompson5910):

oh nvm i see it

jimthompson5910 (jim_thompson5910):

you got x = 2 which is correct

jimthompson5910 (jim_thompson5910):

the graph of `x = 2` is a vertical line because EVERY point on this line has an x coordinate of 2 some points on this line are (2,3), (2,5), (2,7) it doesn't matter what y is. y can be any number you want

OpenStudy (christinestutes):

Ohhhhhhh, ok - I was lost without both variables

OpenStudy (christinestutes):

vertical line on 2 right

jimthompson5910 (jim_thompson5910):

yeah it's just like before when you had that equation with just y in it

OpenStudy (christinestutes):

ok one last thing - if I have equation in slope intercept form, I put y as 0 and solve for x to get x intercept?

jimthompson5910 (jim_thompson5910):

that is correct

OpenStudy (christinestutes):

Thank you so much. I don't need answers but explanations so I can learn this. Many give me just answers and leave me to figure it out on my own!

jimthompson5910 (jim_thompson5910):

I'm glad to be of help.